An aminoacid has pk1-2 pk2-3.5 pk3-6.3
pk4-8 pk5-9.2 pk6-10.4
How to know the isoelectric pH
And how many acidic and basic functional groups are there?
1
Expert's answer
2019-03-01T05:01:12-0500

There are three acidic functional groups which have pK<7 : pK1= 2, pK2 = 3.4, pK3 = 6.3

There are three basic functional gruops whcih have pK>7: pK4 = 8, pK5 = 9.2, pK6 = 10.4

Isoelectric point (pI) is the pH at which an aminoacid is electrically neutral (i.e. exists in the form of zwitterion). To calculate pI we need to average the pKa values relevant to the protonation/deprotonation of the form with no net charge.



The Form IV has net charge = 0, then pI = (pK3 +pK4)/2 = (6.3+8)/2=7.15